OpenStudy (anonymous):

how do i find the greatest common factor of 6,8

OpenStudy (anonymous):

Write down all the factors that they have in common and then choose the biggest one.

OpenStudy (across):

1. Break the numbers down into their prime-power factorization:\[6=2^1\cdot3^1\]\[8=2^3\cdot3^0\]2. Take the minimum of each:\[2^1\cdot3^0=2\]

OpenStudy (eyust707):

thats the fancy way here's how I do it: first I ask myself "what times what equals 6?" 6 times 1 and 2 times 3 "what times what equals 8" 8 times 1 and 2 times 4 "what's the biggest number that both of those have?" 2 =)